OpenStudy (emilyjones284):

Find the vector v having length |v|=8 and direction theta=2pi/3. Write answer in component form.

OpenStudy (ipwnbunnies):

\[x-coordinate = |v|*\cos(\theta)\] \[y = |v|*\sin(\theta)\]

OpenStudy (emilyjones284):

So y=8 times sin(2pi/3)?

OpenStudy (johnweldon1993):

^ Correct

OpenStudy (emilyjones284):

And then y=8 times sin square root of three/2?

OpenStudy (whpalmer4):

\[y = 8\sin(\frac{2\pi}{3}) = 8*\frac{\sqrt{3}}2=\]

OpenStudy (whpalmer4):

\[x= 8\cos(\frac{2\pi}3) = 8*(-\frac{1}{2}) = \]

OpenStudy (emilyjones284):

So 8 square root of three/2 and -4?

OpenStudy (whpalmer4):

You can simplify the first one a bit more, can't you?

OpenStudy (emilyjones284):

would it be \[4\sqrt{3}?\]

OpenStudy (whpalmer4):

Yes.

OpenStudy (emilyjones284):

Oh alright got it... and then how would I put all of that into component form?

OpenStudy (whpalmer4):

Beats me, but I'm sure your book has an example of something in component form, doesn't it? Your two components are \(4\sqrt{3},-4\), that's all I know for sure.

OpenStudy (ipwnbunnies):

It's the x and y component of a coordinate. (x,y)

OpenStudy (whpalmer4):

I'd hate to be responsible for you getting a problem that you've correctly solved marked wrong because I told you the wrong way to format the correct answer...

OpenStudy (ipwnbunnies):

If you graph that point, and the vector at the same angle, it'll be at the same spot.

OpenStudy (emilyjones284):

okay no problem thank you both!

OpenStudy (whpalmer4):

I think, but do not promise, that it is just <x,y>

OpenStudy (whpalmer4):

\[<4\sqrt{3},-4>\]

OpenStudy (emilyjones284):

Okay yeah I think you're right, thanks

OpenStudy (ipwnbunnies):

Oh, right. You're writing the vector, not the point. My bad.